OpenStudy (anonymous):

solve the system of equations by graphing, f(x)=-1/2x +2 g(x)=9/2x -8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

someone please help me

OpenStudy (anonymous):

but its pretty simple

OpenStudy (anonymous):

not for me i really suck at math, please help me step by step

OpenStudy (dumbcow):

is this correct f(x) = -(1/2)x + 2 f(x) = (9/2)x - 8

OpenStudy (anonymous):

not sure how to make the fraction symbols for -1/8 and 9/2

OpenStudy (dumbcow):

thats ok me neither. both of these are linear of the form y = mx+b m = slope (rise/run) b = y-intercept (where it crosses y-axis) Now we need to graph each one First identify y-intercept for our first point 1) b = 2 2) b = -8 put these points on the graph on the y-axis

OpenStudy (dumbcow):

Next we look at slope to find second point 1) m = -1/2 This means starting at our y-intercept we go down 1 and over to right 2, repeat process for additional points 2) m=9/2 This means starting at y-intercept go up 9 and over to right 2 connect points to make each line point where they intersect is the solution

OpenStudy (anonymous):

in the begining the first points b=2 b=-8 is that graphed as (2,-8)?

OpenStudy (dumbcow):

oh sorry i meant 1) to mean line 1 and 2) to mean line 2 they are separate points the x value is always zero for the y-intercept (0,2) (0,-8)

OpenStudy (anonymous):

ok im really bad at this stuff

OpenStudy (dumbcow):

when you think y-intercept, think y-axis....points will be directly on y-axis

OpenStudy (anonymous):

ok, so back to m=-1/2 and m=9/2

OpenStudy (dumbcow):

yes start with the first one go to starting point (0,2) -1 means go down 1 2 means go right 2 plot a point

OpenStudy (anonymous):

ok so after i graphed them i came up with (2,1)?

OpenStudy (dumbcow):

correct this is the point where they cross

OpenStudy (anonymous):

ok do u have time to help me with more

OpenStudy (dumbcow):

for a little bit, if nothing else i'll give you the steps

OpenStudy (anonymous):

ok f(x)=-1/8x-3 g(x)=5/8x-9

OpenStudy (dumbcow):

same thing solve by graphing

OpenStudy (dumbcow):

do the exact same steps we did before the only thing that changes is the slope and y-intercept

OpenStudy (dumbcow):

for f(x) = mx + b m=-1/8 b=-3

OpenStudy (dumbcow):

can you graph the first point from there

OpenStudy (anonymous):

start at 0 and go down the y to -3

OpenStudy (dumbcow):

good

OpenStudy (anonymous):

same for (0,-8)

OpenStudy (anonymous):

they intercept at (2,-3) is that correct?

OpenStudy (dumbcow):

careful the -8 was part of the slope for the first line

OpenStudy (dumbcow):

the other y-intercept would be (0,-9)

OpenStudy (anonymous):

right , after i graphed them i got (2,-3)

OpenStudy (dumbcow):

remember what the difference between m and b are another way to think of it is the slope is always the number associated with the x, y-intercept is always the constant value

OpenStudy (dumbcow):

no you should get (8,-4)

OpenStudy (dumbcow):

starting at (0,-3) go down 1 over right 8

OpenStudy (anonymous):

ok i got it

OpenStudy (dumbcow):

ok good keep repeating those steps gotta go good luck

OpenStudy (anonymous):

thank you so much

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

thank you so much